OpenStudy (anonymous):

multiplying polynomials (5a+6)(a-1)

OpenStudy (anonymous):

5a^2-5a+6a-6=5a^2+a-6

OpenStudy (aroub):

5a^2-5a +6a-6 5a^2+a-6

OpenStudy (anonymous):

FOIL... (first outter inner last) multiply the terms in that order

OpenStudy (anonymous):

5a^2-5a+6a-6 =5a^2+a-6

OpenStudy (anonymous):

\[5a^2 - 5a + 6a - 6 = 5a^2 + a - 6\]

OpenStudy (anonymous):

Use the foil method as mentioned by SJSharks003. FOIL stand for (First-outer-inner-last) (First) Multiply 5a by a to get \[5a^2\]. (Outer) Then multiply 5a by -1 to get -5a. (Inner) Then multiply 6 by a to get 6a. (Last) Then multiply 6 by -1 to get -6. Hope this helps :D

OpenStudy (anonymous):

thnx alot everyone